OpenStudy (anonymous):

muiltiply (5√2+3√3)^2

OpenStudy (anonymous):

I got 30√6-77

OpenStudy (theviper):

\[(a+b)^2=a^2+b^2+2ab\]

OpenStudy (anonymous):

You almost got it right, but I don't get why the 77 is negative though...

OpenStudy (theviper):

Therefore,\[(5\sqrt{2})^2+(3\sqrt{3})^2+2(5\sqrt{2})(3\sqrt{3})\]

OpenStudy (theviper):

Solve it & than u can get it:)

OpenStudy (anonymous):

 your answer is \[77+ 30\sqrt{6}\] hope that helps :)